# Workshop SIG Knowledge Management (FG WM) on 16th September 2025 at KI 2025, Potsdam, Germany

### v2

@@ 2.2 Prompt creation as node classification problem
A major challenge of common RAG approaches is the selection of appropriate text chunks w.r.t a given query, chiefly because of the reliance on embedding-based similarity for textual chunk selection, which can be unreliable or lack coverage due to fluctuating chunk sizes and topic diversity within queries. To this end, our approach provides a natural remedy. Applying a transformation to the constructed knowledge graph, we formulate LLM prompt creation as an unsupervised node classification problem. Thus, we apply a transformation to parts of our tripartite knowledge graph $G$ related to an object of investigation $o$ to create a novel classification graph $G^{\prime}_{o}$ that represents the information in each relationship $e_{c,t}$ between $c$ and $t$ related to $o$ - i.e. given a specific relationship $e_{o,c}$ between $o$ and $c$ - as a node $x\in 0,1$.
Here, $x$ in $G^{\prime}_{o}$ describes the state of the information stored in $e_{c,t}$ to be or not to be included in the prompt related to a query on $o$, i.e. $x=1$ or 0 otherwise. The state of $x$ is governed by the (semantic) relevance of the stored information in $e_{c,t}$ related to $o$’s association with $c$, i.e. $e_{o,c}$ that can be measured as $w_{o,c,t}=cos(E_{e_{o,c}},E_{e_{c,t}})$. Here $E_{e_{o,c}},E_{e_{c,t}}$ denote the vector embeddings of the information stored in $e_{o,c}$ and $e_{c,t}$ with $cos(.)$ denoting the cosine similarity. By collecting all $w_{o,c,t}$ for all $o$ and $t$ connected to a specific concept $c$, we derive an empirical probability distribution $P_{c}$ per each $c$ for all observed similarity scores $w_{o,c,t}$ w.r.t. $c$. Thus, $x$ can be assigned a probability $P_{c}(w_{o,c,t})$ w.r.t. the concept $c$’s specific $P_{c}$.
If $P_{c}(w_{o,c,t})>\alpha$ then $x=1$ with $\alpha$ denoting some hyper-parameter on $P_{c}$, e.g. $\alpha=0.9$. Further, in $G^{\prime}_{o}$ two nodes $x_{i},x_{j}$ are connected, if their respective $e_{c_{i},t}$, $e_{c_{j},t}$ form a first-order connection via a common $t$ in $G$, indicating that text chunk $t$ discusses both concepts $c_{i}$ and $c_{j}$ (see illustration in figure  3). Hence, let’s assume there exists both a connection $e_{o,c_{i}}$ and $e_{o,c_{j}}$ in $G$, while $P_{c_{i}}(w_{o,c_{i},t})>\alpha$ and $P_{c_{j}}(w_{o,c_{j},t})<\alpha$, i.e. $x_{i}=1$ and $x_{j}=0$. Then we can introduce a second, more lenient, condition-specific, interaction-based hyper-parameter $\beta$, e.g. $\beta=0.5$, and evaluate $x_{j}=1$ if $x_{i}=1$ and $P_{c_{j}}(w_{o,c_{j},t})>\beta$ to indicate that the information about $c_{j}$ in $x_{j}$ might still be relevant w.r.t. $o$ - even at a lower absolute $w_{o,c_{j},t}$ - given its co-discussion in $t$ together with $c_{i}$. Note that this step is inspired by principles from Markov Random Fields  [13, 14], which implement a local independence assumption referred to as Markov Blanket  [13].

### v3

@@ 2.2 Prompt creation as node classification problem
A major challenge of common RAG approaches is the selection of appropriate text chunks w.r.t a given query, chiefly because of the reliance on embedding-based similarity for textual chunk selection, which can be unreliable or lack coverage due to fluctuating chunk sizes and topic diversity within queries. To this end, our approach provides a natural remedy. Applying a transformation to the constructed knowledge graph, we formulate LLM prompt creation as an unsupervised node classification problem (see illustration in figure 3). Thus, we apply a transformation to parts of our tripartite knowledge graph $G$ related to an object of investigation $o$ to create a novel classification graph $G^{\prime}_{o}$ that represents the information in each relationship $e_{c,t}$ between $c$ and $t$ related to $o$ - i.e. given a specific relationship $e_{o,c}$ between $o$ and $c$ - as a node $x\in 0,1$.
Here, $x$ in $G^{\prime}_{o}$ describes the state of the information stored in $e_{c,t}$ to be or not to be included in the prompt related to a query on $o$, i.e. $x=1$ or 0 otherwise. The state of $x$ is governed by the (semantic) relevance of the stored information in $e_{c,t}$ related to $o$’s association with $c$, i.e. $e_{o,c}$ that can be measured as $w_{o,c,t}=cos(E_{e_{o,c}},E_{e_{c,t}})$. Here $E_{e_{o,c}},E_{e_{c,t}}$ denote the vector embeddings of the information stored in $e_{o,c}$ and $e_{c,t}$ with $cos(.)$ denoting the cosine similarity. By collecting all $w_{o,c,t}$ for all $o$ and $t$ connected to a specific concept $c$, we derive an empirical probability distribution $P_{c}$ per each $c$ for all observed similarity scores $w_{o,c,t}$ w.r.t. $c$. Thus, $x$ can be assigned a probability $P_{c}(w_{o,c,t})$ w.r.t. the concept $c$’s specific $P_{c}$. If $P_{c}(w_{o,c,t})>\alpha$ then $x=1$ with $\alpha$ denoting some hyper-parameter on $P_{c}$, e.g. $\alpha=0.9$.
@@
See figure 4 for an example. $P_{c}$ is derived based on collecting all six $w_{o,c,t}=cos(E_{e_{o,c}},E_{e_{c,t}})$ for all $o\in$ (Peter Mueller, Barbara Schmidt, Oskar Lehm) and $t\in$ (text chunk I, text chunk II). E.g. let us assume $w_{o=\text{Barbara Schmidt},c,t={\text{text chunk II}}}=cos(E_{e_{o=\text{Barbara Schmidt},c}},E_{e_{c,t={\text{text chunk II}}}})=0.7$ with $c=\text{blood pressure}$, hence forming the top element of the empirical distribution (illustrated as histogram in figure 4). For the given example, $P_{c}(w_{o={\text{Barbara Schmidt}},c,t={\text{text chunk II}}})>\alpha$, therefore, if the provided query would be related to the cross-examination of object $o={\text{'Barbara Schmidt'}}$, this would result in the selection of $t={\text{'text chunk II'}}$ as context for the prompt to answer the query. To this end, this fast, i.e. embedding based, pre-evaluation of potentially relevant text chunks based on applying some hyper-parameter $\alpha$ on $P_{c}$ not only serves as a pre-filter in order to avoid cluttering the language model context window but also as a measure to facilite scalability as the number of objects $o$ and text elements $t$ related to a concept $c$ increases.
Figure 4: Illustration of the derivation of $P_{c}$ via collection of all combinations $w_{o,c,t}=cos(E_{e_{o,c}},E_{e_{c,t}})$ for all $o\in$ (Peter Mueller, Barbara Schmidt, Oskar Lehm) and $t\in$ (text chunk I, text chunk II), as illustrated in the respective table and exemplary histogram. E.g. assuming $w_{o={\text{Barbara Schmidt}},c,t={\text{text chunk II}}}=cos(E_{e_{o={\text{Barbara Schmidt}},c}},E_{e_{c,t={\text{text chunk II}}}})=0.7$ with $c=\text{'blood pressure'}$, therefore, $P_{c}(w_{o={\text{Barbara Schmidt}},c,t={\text{text chunk II}}})>\alpha$. Given the provided query to be related to object $o={\text{'Barbara Schmidt'}}$ would result in the selection of $t={\text{'text chunk II'}}$.
Further, in $G^{\prime}_{o}$ two nodes $x_{i},x_{j}$ are connected, if their respective $e_{c_{i},t}$, $e_{c_{j},t}$ form a first-order connection via a common $t$ in $G$, indicating that text chunk $t$ discusses both concepts $c_{i}$ and $c_{j}$ (see illustration in figure 3). Hence, let’s assume there exists both a connection $e_{o,c_{i}}$ and $e_{o,c_{j}}$ in $G$, while $P_{c_{i}}(w_{o,c_{i},t})>\alpha$ and $P_{c_{j}}(w_{o,c_{j},t})<\alpha$, i.e. $x_{i}=1$ and $x_{j}=0$. Then we can introduce a second, more lenient, condition-specific, interaction-based hyper-parameter $\beta$, e.g. $\beta=0.5$, and evaluate $x_{j}=1$ if $x_{i}=1$ and $P_{c_{j}}(w_{o,c_{j},t})>\beta$ to indicate that the information about $c_{j}$ in $x_{j}$ might still be relevant w.r.t. $o$ - even at a lower absolute $w_{o,c_{j},t}$ - given its co-discussion in $t$ together with $c_{i}$. Note that this step is inspired by principles from Markov Random Fields  [14, 15], which implement a local independence assumption referred to as Markov Blanket  [14].